### v1

# Simulation Based Bayesian Optimization

Abstract
Bayesian Optimization (BO) is a powerful method for optimizing black-box functions by combining prior knowledge with ongoing function evaluations. BO constructs a probabilistic surrogate model of the objective function given the covariates, which is in turn used to inform the selection of future evaluation points through an acquisition function. For smooth continuous search spaces, Gaussian Processes (GPs) are commonly used as the surrogate model as they offer analytical access to posterior predictive distributions, thus facilitating the computation and optimization of acquisition functions. However, in complex scenarios involving optimizations over categorical or mixed covariate spaces, GPs may not be ideal.
This paper introduces Simulation Based Bayesian Optimization (SBBO) as a novel approach to optimizing acquisition functions that only requires sampling-based access to posterior predictive distributions. SBBO allows the use of surrogate probabilistic models tailored for combinatorial spaces with discrete variables. Any Bayesian model in which posterior inference is carried out through Markov chain Monte Carlo can be selected as the surrogate model in SBBO. In applications involving combinatorial optimization, we demonstrate empirically the effectiveness of SBBO method using various choices of surrogate models.
Keywords: Acquisition functions, Bayesian decision theory, Black-box functions, Combinatorial optimization, Markov chain Monte Carlo
1 Introduction
The problem of optimizing a potentially nonlinear function $f(\boldsymbol{x})$ over a certain set $\mathcal{S}$ has been extensively studied. In many applications, the objective function is not explicitly known. This can make optimization challenging, as it is difficult to determine the optimal solution without accurate knowledge of the function. Moreover, $f(\boldsymbol{x})$ may be expensive to evaluate and a budget constraint often restricts evaluations of the objective function to a sparse selection of points, further complicating the optimization task. As an example, consider the process of drug discovery. During its early stages, the goal is to identify candidate molecules with a desired property, such as maximum binding affinity to a given biological target. However, measuring this binding affinity requires synthesizing and testing molecules in vitro, limiting the number of leads that can be evaluated. Furthermore, expressing affinity as a mathematical function of the molecule’s structure is not possible.
Bayesian Optimization (BO, Mockus (1994)) stands as an efficient approach for identifying the extrema of a black-box objective function. Its effectiveness originates from its capacity to integrate prior knowledge about the function with the evaluations previously conducted, enabling informed decisions on where to carry out additional evaluations. Bayesian Optimization involves the construction of a Bayesian surrogate model for $f(\boldsymbol{x})$. Subsequently, the next evaluation point is selected through the maximization of an acquisition function (Frazier, 2018). Commonly used acquisition functions, including expected improvement, are defined as the expected values of utility functions computed with respect to the posterior predictive distribution of $f(\boldsymbol{x})$ derived from the Bayesian surrogate model. In the remainder of the paper, we will restrict our attention to this class of acquisition functions.
Much of the BO literature focuses on continuous search spaces, where Gaussian Processes (GPs, Rasmussen et al. (2006)) are predominantly chosen as the Bayesian surrogate model. GPs are well-suited to optimization tasks in continuous search spaces as they rely on the smoothness defined by a kernel to model uncertainty. Moreover, GPs offer the advantage of yielding a analytical posterior predictive distribution (the posterior is Gaussian), enabling the analytical computation of commonly used acquisition functions like expected improvement or probability of improvement. However, in more intricate scenarios, such as those involving combinatorial (i.e. categorical) or mixed categorical-continuous covariate spaces (i.e. where the search space is a combination of discrete and continuous variables), GPs may not be the most suitable choice.
Related work.
In the last few years there has been some research in how to perform Bayesian optimization over combinatorial domains. As highlighted by Wang et al. (2023), this task comes with two major challenges: the construction of effective surrogate models over the combinatorial space, and the effective search in the combinatorial domain for the next structure for evaluation according to the acquisition function. Most previous research has been devoted to the first task. Within this, an important line of investigation consists of modifying GP kernel functions to accommodate categorical covariate spaces. For instance Garrido-Merchán and Hernández-Lobato (2020) treat categorical variables as continuous via one-hot encoding. Then, within the kernel function, continuous covariates are suitably transformed to one-hot encodings. Oh et al. (2019) constructs a combinatorial graph over the combinatorial search space. Subsequently, graph Fourier transforms are used to derive a diffusion kernel on the graph.
Another interesting line of research performs BO over a continuous latent representation of the combinatorial spaces learned through a Variational Autoencoder (Gómez-Bombarelli et al., 2018). Finally, a third line of research gets rid of the GPs, and proposes using inherently discrete surrogate models. For instance Hutter et al. (2011) use random forests to model the objective function given the covariates. Baptista and Poloczek (2018) model the response as a linear function of the covariates including all possible pairwise interactions. A heavy-tailed horseshoe prior is used to induce sparsity on the coefficients.
In this paper, we focus on the second challenge: the task of effective search within the combinatorial domain. We introduce Simulation Based Bayesian Optimization (SBBO), a novel approach to optimize acquisition functions in Bayesian Optimization. Unlike previous approaches, SBBO does not require knowing the posterior predictive distribution analytically. All that is needed is sampling-based access to posterior predictive distributions. This allows us to use any Bayesian model whose posterior inference is done via Markov chain Monte Carlo (MCMC). Thus, determining where to make the next function evaluation becomes a simpler task that does not rely on closed-form results that are typically only available under the Gaussian surrogate. SBBO opens the door to using surrogate probabilistic models that may be better tailored for Bayesian optimization tasks in combinatorial spaces, where the search for optimal solutions becomes more challenging due to the presence of discrete variables.
The paper is structured as follows: Section 2 provides an overview of Bayesian Optimization. Section 3 presents SBBO and summarises the surrogate probabilistic models that we use for the experiments presented in Section 4. Finally, Section 5 gives some concluding remarks and lines of future research.
2 Bayesian Optimization
Bayesian optimization (BO) is a class of methods centered on solving the optimization problem
$\boldsymbol{x}^{*}=\max_{\boldsymbol{x}\in\mathcal{S}}f(\boldsymbol{x}),$ (2.1)
In the setting of Bayesian optimization, the objective function $f(\boldsymbol{x})$ is assumed to be a black box, so no assumptions of convexity or linearity may be made. Additionally, it is assumed that $f$ is expensive to evaluate in that the number of evaluations of $f$ that can be feasibly performed is small, perhaps due to computing time or financial cost. It is further assumed that querying the function at $x$ provides a noisy observation $y$ of $f(x)$, with normal zero mean, $\sigma^{2}$ variance noise.
BO uses a surrogate probabilistic model of the objective function, usually obtained through Bayesian methods, to inform the next evaluation decision. It works in an iterative algorithm by 1) fitting the surrogate model and 2) using an acquisition function to obtain a new point $\boldsymbol{x}$ at which we evaluate or observe $f(\boldsymbol{x})$. Specifically, the surrogate model is fitted using all available data $\mathcal{D}_{1:n}=\{(\boldsymbol{x}_{i},y_{i}),i=1,\ldots,n\}$. Then using an acquisition function, we select a new $\boldsymbol{x}_{n+1}$ and obtain a corresponding $y_{n+1}$. The surrogate model is then updated to include the additional data point $(\boldsymbol{x}_{n+1},y_{n+1})$. We briefly review each of these two key components of Bayesian Optimization: the surrogate model and the acquisition function.
The following observational model is assumed
[…]
where $f(\boldsymbol{x}_{t})$ could be parametric or non-parametric. Given a new location $\boldsymbol{x}$ and appropriate choice of prior, the posterior predictive distribution on $f(\boldsymbol{x}_{t})$ is taken as the surrogate model. If a parametric model is chosen for $f$, then $f(\boldsymbol{x}_{t})$ will depend on parameters $\boldsymbol{\beta}$ and is denoted as $f_{\boldsymbol{\beta}}(\boldsymbol{x}_{t})$. A prior $\pi(\boldsymbol{\beta})$ for the parameters $\boldsymbol{\beta}$ needs to be specified. Coupled with a likelihood $\pi(y|\boldsymbol{x},\boldsymbol{\beta},\sigma^{2})$, after observing data $\mathcal{D}_{1:n}$, we obtain the posterior $\pi(\boldsymbol{\beta}|\mathcal{D}_{1:n})$. Then, the surrogate model is
$\pi(f(\boldsymbol{x})|\boldsymbol{x},\mathcal{D}_{1:n})\equiv\pi(f_{\boldsymbol{\beta}}(\boldsymbol{x})|\boldsymbol{x},\mathcal{D}_{1:n})=\mathbb{E}_{\boldsymbol{\beta}|\mathcal{D}_{1:n}}\left[f_{\boldsymbol{\beta}}(\boldsymbol{x})\right]$
If we choose instead a non-parametric model for $f$, then a prior $\pi(f)$ over $f$ needs to be specified. Coupled with the likelihood $\pi(y|f,\sigma^{2})$, we obtain the posterior distribution $\pi(f|\mathcal{D}_{1:n})$ over $f$. Then for a given new location $\boldsymbol{x}$, the evaluation of the posterior at $\boldsymbol{x}$, $\pi(f(\boldsymbol{x})|\boldsymbol{x},\mathcal{D}_{1:n})$, is taken as the surrogate model.
An acquisition function is used to direct the search for the optimal $\boldsymbol{x}^{*}\in\mathcal{S}$. There are several possible choices. In this paper we restrict our attention to the class of acquisition functions that can be written as expectations of a utility function with respect to the posterior $\pi(f(\boldsymbol{x})|\boldsymbol{x},\mathcal{D}_{1:n})$. Some of the most commonly used acquisition functions belong to this class, such as expected improvement and probability of improvement. Let $u(f(\boldsymbol{x}),\boldsymbol{x})$ denote the utility achieved by evaluating the function at a new point $\boldsymbol{x}$ and obtaining $f(\boldsymbol{x})$.
Thus, the next point $\boldsymbol{x}_{n+1}$ is chosen as the one that maximizes the acquisition function, i.e. the expectation of the utility function:
$\boldsymbol{x}_{n+1}=\arg\max_{\boldsymbol{x}\in\mathcal{S}}\mathbb{E}_{f|\boldsymbol{x},\mathcal{D}_{1:n}}\left[u(f(\boldsymbol{x}),\boldsymbol{x})\right].$ (2.3)
Under the choice of GP prior for $f$, the posterior is Gaussian and under common choices of the utility functions, there is a closed-form expression for the expectation (2.3). The BO framework is indeed flexible, but computationally efficient and analytic results rely on this choice of Gaussian process prior. However, in specific scenarios-such as when the search space comprises categorical or mixed categorical-continuous variables—this GP prior may not be the best fit. In these situations, various alternative Bayesian models better tailored to the problem may be available. However, the alternatives often lack an analytically-defined posterior distribution, obstructing the selection of the next point $\boldsymbol{x}_{n+1}$ according to (2.3). We note, however, that we can sample from the posterior. To address this challenge presented by a non-GP choice of prior, we introduce a Bayesian optimization methodology that exclusively requires sampling access to the posterior, enabling the use of Bayesian models better suited to structured covariates spaces.
It is important to emphasize that choosing the next sampling location in (2.3) implicitly entails adopting a “one-step” assumption: the selection is made under the presumption that no future evaluations will take place. When there is budget for more than one function evaluation, Expected Improvement optimization unequivocally adopts a myopic approach. While several alternative strategies have been proposed for multi-step lookahead Bayesian Optimization (Wu and Frazier, 2019; Lam et al., 2016), we focus on the one-step look-ahead case.
3 Simulation Based Bayesian Optimization
We incorporate ideas from the simulation-based optimal Bayesian design literature (Müller, 2005) into the field of Bayesian Optimization, yielding our proposed method of Simulation-Based Bayesian Optimization (SBBO).
3.1 Main Idea
As in all Bayesian Optimization tasks, we aim to solve the optimization problem 2.1. The goal is to estimate the objective function $f$ with few evaluations. We do not assume that the objective function $f$ is continuous, but otherwise all the difficulties of estimating $f$ hold.
Imagine we have access to (perhaps through previous experiments or function evaluations) data $\mathcal{D}_{1:n}=\{(\boldsymbol{x}_{i},y_{i}),i=1,\ldots,n\}$, where $\boldsymbol{x}_{i}$ is a vector of covariates for observation $i$ and $y_{i}$ is obtained from the observation model in (2.2). Suppose that we have a budget of at most $M$ more function evaluations. Using all of the available information, the task is to determine the next $\boldsymbol{x}_{n+1},\boldsymbol{x}_{n+2},\ldots,\boldsymbol{x}_{n+M}$ at which we will evaluate the objective function. As is common practice in BO, rather than deciding all $M$ new locations/vectors at once, we proceed in a greedy manner by deciding on new locations in an iterative method.
To select the next point, $\boldsymbol{x}_{n+1}$, we typically maximize an acquisition function computed using $\pi(f(\boldsymbol{x})|\boldsymbol{x},\mathcal{D}_{1:n})$ over possible values of $\boldsymbol{x}$. Let $u(f(\boldsymbol{x}),\boldsymbol{x})$ be a non-negative and bounded utility function that quantifies the benefits/utility/gains from evaluating the objective function at $\boldsymbol{x}$ and observing the corresponding $f(\boldsymbol{x})$, assuming we have already observed data $\mathcal{D}_{1:n}$. Following Bayesian decision-theoretic principles, the next evaluation should be performed at the location $\boldsymbol{x}$ that maximizes the expected utility
[…]
In order to perform Bayesian optimization, all we need is to (1) decide on a suitable Bayesian surrogate model and (2) solve (3.1). As previously mentioned, except in special cases, the posterior $\pi(f(\boldsymbol{x})|\boldsymbol{x},\mathcal{D}_{1:n})$ is typically only accessible through sampling. For this purpose, we borrow ideas from the optimal Bayesian design community. The key idea is to reframe optimization of (3.1) as a simulation problem. Following Bielza et al. (1999), we define an augmented distribution $g$ that incorporates $\boldsymbol{x}$ as a random variable:
$g(\boldsymbol{x},f(\boldsymbol{x})|\mathcal{D}_{1:n})\propto u(f(\boldsymbol{x}),\boldsymbol{x})\pi(f(\boldsymbol{x})|\boldsymbol{x},\mathcal{D}_{1:n}).$ (3.2)
It is straightforward to see that the marginal of $g$ in $\boldsymbol{x}$ is proportional to the expected utility defined in (3.1). Thus, the mode of this marginal distribution coincides with the optimal next evaluation point. This suggests a simulation-based approach to find $\boldsymbol{x}_{n+1}$: we can simply simulate from $(\boldsymbol{x},f(\boldsymbol{x}))\sim g(\boldsymbol{x},f(\boldsymbol{x})|\mathcal{D}_{1:n})$ using standard MCMC techniques and compute the mode of $g(\boldsymbol{x}|\mathcal{D}_{1:n})$ using the generated samples $\boldsymbol{x}$. However, this strategy is limited to low-dimensional $\boldsymbol{x}$. For high-dimensional covariate spaces, estimating the mode of $g(\boldsymbol{x}|\mathcal{D}_{1:n})$ from the simulation output becomes impracticable. Müller et al. (2004) propose an alternative augmented distribution $g_{H}$ defined as
$g_{H}(\boldsymbol{x},f(\boldsymbol{x})_{1},\ldots,f(\boldsymbol{x})_{H}|\mathcal{D}_{1:n})\propto\prod_{h=1}^{H}u(f(\boldsymbol{x})_{h},\boldsymbol{x})\pi(f(\boldsymbol{x})_{h}|\boldsymbol{x},\mathcal{D}_{1:n}),$ (3.3)
where $f(\boldsymbol{x})_{1},\ldots,f(\boldsymbol{x})_{H})$ are $H$ copies of the random variable $f(\boldsymbol{x})$. Note that the marginal of $g_{H}$ in $\boldsymbol{x}$ is
$g_{H}(\boldsymbol{x}|\mathcal{D}_{1:n})\propto\Psi(\boldsymbol{x})^{H}.$ (3.4)
Thus, the mode of this marginal distribution also coincides with the optimal point at which we should make our next function evaluation. In particular, if $H$ is sufficiently large, the marginal distribution $g_{H}(\boldsymbol{x})$ is tightly concentrated on the next $\boldsymbol{x}_{n+1}$.
Müller et al. (2004) propose an inhomogeneous MCMC sampling scheme to simulate from $g_{H}$. They begin by generating samples from $g$ using the standard Metropolis-Hastings algorithm. In subsequent iterations, the target distribution is changed to $g_{H}$, with $H$ increasing across iterations. The inhomogeneous MCMC is constructed in such a way that the stationary distribution for fixed $H$ is $g_{H}$. We use this idea to find the next evaluation point in the context of Bayesian optimization. The algorithm is illustrated in Algorithm 1, where a Metropolis-Hastings sampling scheme is used.
Algorithm 1       Simulation Based Bayesian Optimization
Obtain initial dataset $\mathcal{D}_{1:n}=\{(\boldsymbol{x}_{i},y_{i}),i=1,\ldots,n\}$. Determine form of utility function $u_{\mathcal{D}}(\boldsymbol{x},y)$.
Perform SBBO to find the next evaluation location.
  1. 1.
    Choose initial $\boldsymbol{x}_{n+1}^{(0)}$. For $h=1,\ldots,H$, simulate $f(\boldsymbol{x})^{(0)}_{h}\sim\pi(f(\boldsymbol{x})|\boldsymbol{x}_{n+1}^{(0)},\mathbf{D}_{1:n})$ and evaluate corresponding utility $u^{(0)}_{h}=u(f(\boldsymbol{x}_{n+1}^{(0)}),\boldsymbol{x}_{n+1}^{(0)})$. Define $v^{(0)}=\frac{1}{H}\sum_{h=1}^{H}\log u^{(0)}_{h}$.
  2. 2.
    Generate candidate $\tilde{\boldsymbol{x}}_{n+1}$ from proposal distribution $q(\tilde{\boldsymbol{x}}_{n+1}|\boldsymbol{x}_{n+1}^{(0)})$.
  3. 3.
    For $h=1,\ldots,H$, simulate $\tilde{f}(\boldsymbol{x})_{h}\sim\pi(f(\boldsymbol{x})|\tilde{\boldsymbol{x}}_{n+1},\mathcal{D}_{1:n})$ and evaluate corresponding utility $\tilde{u}_{h}=u(\tilde{f}(\boldsymbol{x})_{h},\tilde{\boldsymbol{x}}_{n+1})$. Define $\tilde{v}=\frac{1}{H}\sum_{h=1}^{H}\log\tilde{u}_{h}$.
  4. 4.
    Compute
    $\alpha$ $=\min\left\{1,\exp(H\tilde{v}-Hv^{(0)})\frac{q(\boldsymbol{x}^{(0)}_{n+1}|\tilde{\boldsymbol{x}}_{n+1})}{q(\tilde{\boldsymbol{x}}_{n+1}|\boldsymbol{x}^{(0)}_{n+1})}\right\}$
  5. 5.
    Set
[…]
  6. 6.
    Increase $H$ according to cooling schedule.
  7. 7.
    Repeat (2) - (6) until convergence.
Select $\boldsymbol{x}_{n+1}$ as the modal value of the simulated $\{\boldsymbol{x}_{n+1}^{(B)},\boldsymbol{x}_{n+1}^{(B+1)},\boldsymbol{x}_{n+1}^{(B+2)},\ldots\}$, where the first $B$ simulated values are removed for burn-in.
Evaluate the true objective at $\boldsymbol{x}_{n+1}$ to obtain $y_{n+1}$.
Update dataset $\mathcal{D}_{1:(n+1)}=\mathcal{D}_{1:n}\cup(\boldsymbol{x}_{n+1},y_{n+1})$
Set $n=n+1$.
If the utility is non-negative and bounded and the state space of the inhomogeneous Markov chain produced by this sampler is discrete, then Müller et al. (2004) show that the chain is strongly ergodic if the updating schedule for $H$ is logarithmic. Moreover, its stationary distribution is uniform over the set of $\boldsymbol{x}$ maximizing (3.1).
A key point of this approach is that once a candidate $\tilde{\boldsymbol{x}}_{n+1}$ has been proposed, the distribution $\pi(f(\boldsymbol{x})|\tilde{\boldsymbol{x}}_{n+1},\mathcal{D}_{1:n})$ is used as proposal for generating $\tilde{f}(\boldsymbol{x})_{h}$ for $h=1,\dots,H$. As a consequence, within each iteration of Algorithm 1, the acceptance ratio is:
$\alpha$ $=$ $\min\left\{1,\frac{\prod_{h=1}^{H}u(\tilde{f}(\boldsymbol{x})_{h},\tilde{\boldsymbol{x}}_{n+1})\pi(\tilde{f}(\boldsymbol{x})_{h}|\tilde{\boldsymbol{x}}_{n+1},\mathcal{D}_{1:n})\pi(f(\boldsymbol{x})^{(0)}_{h}|\boldsymbol{x}^{(0)}_{n+1},\mathcal{D}_{1:n})q(\boldsymbol{x}^{(0)}_{n+1}|\tilde{\boldsymbol{x}}_{n+1})}{\prod_{h=1}^{H}u(f(\boldsymbol{x})^{(0)}_{h},\boldsymbol{x}^{(0)}_{n+1})\pi(f(\boldsymbol{x})^{(0)}_{h}|\boldsymbol{x}^{(0)}_{n+1},\mathcal{D}_{1:n})\pi(\tilde{f}(\boldsymbol{x})_{h}|\tilde{\boldsymbol{x}}_{n+1},\mathcal{D}_{1:n})q(\tilde{\boldsymbol{x}}_{n+1}|\boldsymbol{x}^{(0)}_{n+1})}\right\}$
$=$ $\min\left\{1,\frac{\prod_{h=1}^{H}u(\tilde{f}(\boldsymbol{x})_{h},\tilde{\boldsymbol{x}}_{n+1})q(\boldsymbol{x}^{(0)}_{n+1}|\tilde{\boldsymbol{x}}_{n+1})}{\prod_{h=1}^{H}u(f(\boldsymbol{x})^{(0)}_{h},\boldsymbol{x}^{(0)}_{n+1})q(\tilde{\boldsymbol{x}}_{n+1}|\boldsymbol{x}^{(0)}_{n+1})}\right\}$
$=$ $\min\left\{1,\exp(H\tilde{v}-Hv^{(0)})\frac{q(\boldsymbol{x}^{(0)}_{n+1}|\tilde{\boldsymbol{x}}_{n+1})}{q(\tilde{\boldsymbol{x}}_{n+1}|\boldsymbol{x}^{(0)}_{n+1})}\right\}.$
Thus, the posterior predictive density for $\tilde{f}(\boldsymbol{x})$ cancels out with the proposal density and we do not need to evaluate it. This allows us to optimize (3.1) when just having sampling access to the posterior, thus achieving our goal.
3.2 Gibbs Sampler
A possible variation of Algorithm (1) entails using a Gibbs sampler to simulate from $g_{H}$ rather than a Metropolis-Hastings scheme for every $H$. Recall that $\boldsymbol{x}$ is a vector of $p$ covariates. The conditional distribution of the $s$-th element of $\boldsymbol{x}$ can be written as
$\begin{split}&g_{H}(\boldsymbol{x}_{s}|\boldsymbol{x}_{-s},f(\boldsymbol{x})_{1},\ldots,f(\boldsymbol{x})_{H}),\mathcal{D}_{1:n})\propto\\ &\quad\exp\Bigg{\{}\sum_{h=1}^{H}\log\left[u(f(\boldsymbol{x})_{h},\boldsymbol{x}_{s}\cup\boldsymbol{x}_{-s}\right]+\log\left[\pi(f(\boldsymbol{x})_{h}|\boldsymbol{x}_{s}\cup\boldsymbol{x}_{-s},\mathcal{D}_{1:n})\right]\Bigg{\}}\end{split}$ (3.5)
For the particular case in which $\boldsymbol{x}$ is a vector of $p$ categorical variables, this marginal is simply the softmax distribution over $\sum_{h=1}^{H}\log\left[u(f(\boldsymbol{x})_{h},\boldsymbol{x}_{s}\cup\boldsymbol{x}_{-s}\right]+\log\left[\pi(f(\boldsymbol{x})_{h}|\boldsymbol{x}_{s}\cup\boldsymbol{x}_{-s},\mathcal{D}_{1:n})\right]$. However, sampling from this distribution requires evaluating posterior predictive probabilities, which defeats the purpose of SBBO. As an alternative, we could create a Gibbs sampling scheme to sample from $g_{H}(\boldsymbol{x}_{1},\boldsymbol{x}_{2},\dots,\boldsymbol{x}_{H}|\mathcal{D}_{1:n})$ where $(f(\boldsymbol{x})_{1},\ldots,f(\boldsymbol{x})_{H})$ has been integrated out. Sampling from $g_{H}(\boldsymbol{x}_{s}|\boldsymbol{x}_{-s})$ can be done using Metropolis steps. Assuming that the current state of the chain is $\boldsymbol{x}$, $f(\boldsymbol{x})_{1},\dots,f(\boldsymbol{x})_{H}$ and $v=\frac{1}{H}\sum{h=1}^{H}\log u(f(\boldsymbol{x})_{h},\boldsymbol{x})$, we can proceed as follows:
  1. 1.
    Propose a candidate $\tilde{\boldsymbol{x}}_{s}$ using a probing distribution $q_{s}(\tilde{\boldsymbol{x}}|\boldsymbol{x})$ that only changes the $s$-th element of $\boldsymbol{x}$.
  2. 2.
    For $h=1,\ldots,H$, simulate $\tilde{f}(\boldsymbol{x})_{h}\sim\pi(f(\boldsymbol{x})|\tilde{\boldsymbol{x}}_{s}\cup\boldsymbol{x}_{-s},\mathcal{D}_{1:n})$.
  3. 3.
    Compute $\tilde{v}=\frac{1}{H}\sum_{h=1}^{H}\log u(f(\tilde{\boldsymbol{x}}_{s}\cup\boldsymbol{x}_{-s})_{h},\tilde{\boldsymbol{x}}_{s}\cup\boldsymbol{x}_{-s})$
  4. 4.
    Evaluate acceptance probability
    $\alpha$ $=\min\left\{1,\exp(H\tilde{v}-Hv)\frac{q(\boldsymbol{x}|\tilde{\boldsymbol{x}}_{s}\cup\boldsymbol{x}_{-s})}{q(\tilde{\boldsymbol{x}}_{s}\cup\boldsymbol{x}_{-s}|\boldsymbol{x})}\right\}$
This Gibbs sampling scheme would be embedded in the inhomogenous Markov chain defined in Algorithm 1.
SBBO allows us to naturally accomodate the scenario of mixed categorical-continuous covariate spaces. We simply adapt the proposal distribution of each particular element. For instance, if $\boldsymbol{x}_{q}$ is continuous we could use a Gaussian proposal distribution, while if it is categorical we could use a discrete uniform proposal over all possible choices for the covariate.
3.3 Probabilistic Surrogate Models
We can now optimize (3.1) for any probabilistic surrogate model, as long as we have access to posterior samples of $f$. This flexibility allows us to consider a wide range of surrogate models. In particular, we have tested the following models, which may be well-suited for cases where $\mathcal{S}$ is the space of $p$-dimensional categorical covariates.
Tanimoto Gaussian Process.
This surrogate model is particularly useful when $x\in\{0,1\}^{p}$. The uncertainty on $f(x)$ is modeled through a Gaussian Process with the following kernel function:
$k(x,x^{\prime})=\frac{\phi(x\cdot x^{\prime})}{\|x\|^{2}+\|x^{\prime}\|^{2}-x\cdot x^{\prime}}.$
The numerator is proportional to the number of dimensions where both $x$ and $x^{\prime}$ are 1, and the denominator signifies the total number of dimensions where either $x$ or $x^{\prime}$ or both is 1. The parameter $\phi$ is usually set to the value maximizing the marginal likelihood thus following a empirical Bayes approach. For categorical covariates with more than two possible values, we would need to first convert the covariates into appropriate binary variables (e.g. via one-hot encoding) before computing the previous kernel. Under this model, the posterior predictive distribution is Gaussian with a certain mean and variance analytically available as in any GP.
Sparse Bayesian linear regression.
The response variable is modeled as a linear function of the covariates, including all possible pairwise interactions. A heavy-tailed horseshoe prior is used to induce sparsity.
$y=f_{\boldsymbol{\alpha}}(\boldsymbol{x})+\epsilon=\alpha_{0}+\sum_{j}\alpha_{j}x_{j}+\sum_{i,j>i}\alpha_{ij}x_{i}x_{j}+\epsilon$ (3.6)
$\epsilon\sim\mathcal{N}(0,\sigma^{2})$ (3.7)
$\alpha_{k}|\beta_{k},\tau,\sigma^{2}\sim\mathcal{N}(0,\beta_{k}^{2}\tau^{2}\sigma^{2})$
$\beta_{k},\tau\sim\text{Cauchy}_{[0,\infty)}(1)$
$p(\sigma^{2})\propto\sigma^{-2}$
This model was first proposed by Baptista and Poloczek (2018) in the context of Bayesian Optimization. Posterior inference is achieved through Gibbs sampling, as detailed in Baptista and Poloczek (2018). This method allows us to sample from the posterior of $\alpha_{k},\beta_{k},\tau$, and $\sigma^{2}$, facilitating easy draws from the posterior of $f_{\alpha}(\boldsymbol{x})$ that could be directly utilized in Algorithm 1. In this scenario, access to analytical posterior predictive distributions is unavailable, making the computation of common acquisition functions challenging. To address this, Baptista and Poloczek (2018) propose a tailored acquisition function: a single posterior sample of $\alpha$ is produced, and the acquisition function is defined as $f_{\alpha}(\boldsymbol{x})-\lambda\|\boldsymbol{x}\|_{1}$, where a penalty term controlled by $\lambda$ is included. This acquisition function now has a closed-form, appearing as a quadratic form. Its optimization can be reformulated as a semidefinite program that can be efficiently approximated in polynomial time to a desired precision. With SBBO, there is no need for tailored acquisition functions as we just require having sampling access to the posterior predictive distribution. This enables using model (3.6) with the common acquistion functions in the BO literature.
NGBoost.
Natural Gradient Boosting (NGBoost, Duan et al. (2020)) is a non-Bayesian algorithm for generic probabilistic prediction via gradient boosting. In NGBoost, the conditional distribution of the output given covariates is modelled through a parametric distribution $y|\boldsymbol{x}\sim P_{\theta}(\boldsymbol{x})$. The parameters $\theta(\boldsymbol{x})$ are obtained through an additive combination of $M$ base learners and an initial $\theta^{(0)}$.
$\theta=\theta^{(0)}-\eta\sum_{m=1}^{M}\rho^{(m)}\cdot f^{(m)}(x)$
where $\rho^{(m)}$ are scaling factors and $\eta$ is a common learning rate. The base learners are trained to minimize a proper scoring rule using a refinement of the gradient boosting algorithm. For this training process, NGBoost employs the natural gradient, a variant of gradient descent that takes into account the underlying geometry of probability distributions. This leads to faster convergence and improved stability during training.
A notable strength of NGBoost lies in its adaptability, as it can be paired with various base learners. Our experiments focus on two such learners: shallow decision trees (referred to as NGBoost-dec) and linear regressions with lasso regularization (named NGBoost-linCV). For both these configurations, we utilize the log-score as the scoring rule.
Bayesian Neural Network.
Bayesian Neural Networks (BNNs) are probabilistic deep learning models that replace deterministic weight values of a neural network with weight distributions. This allows the model to capture the existing epistemic uncertainty in the predictions due to limited training data. In our experiments, we employ a simple three- hidden layers fully-connected BNN whose inputs are the one-hot encoded covariates. We train this model using variational inference, combined with the reparametrization estimator technique (Kingma and Welling, 2013).
4 Experiments
We conduct experiments in three different combinatorial optimization problems [^1] All code to reproduce the experiments in this paper has been open sourced and is available at https://github.com/roinaveiro/sbbo.: a binary quadratic program with 10 variables, a contamination control in a food supply chain with 25 stages, and RNA sequence optimization with a sequence length of 30. Details of the problem settings are provided in the subsequent subsections. We evaluate the performance of SBBO with the various probabilistic surrogate models described at the end of the previous section: Sparse Bayesian linear regression with interactions (sbbo-BLr), Bayesian Neural Network (sbbo-BNN), Tanimoto Gaussian Process Regression (sbbo-GPr), natural gradient boosting with a shallow decision tree model as base learner (sbbo-NGBdec), and natural gradient boosting with an sparse linear regression as base learner (sbbo-NGBlinCV).
We compare performance of these SBBO approaches with two benchmarks: simulated annealing (SA), a widely used algorithm in combinatorial optimization problems (Spears, 1993) and random local search (RS, Bergstra and Bengio (2012)).
The experimental setup was design as follows. We start with a random initial dataset containing five pairs of covariates and responses. For each probabilistic surrogate model and fixed number of iterations $t$, we follow these steps:
  1. 1.
    Use the surrogate model to determine the next evaluation point, utilizing the expected improvement as the acquisition function. Thus, the utility function is defined as $u(\boldsymbol{x})=\max\left[f(\boldsymbol{x})-f^{*}(\boldsymbol{x}),0\right]$, where $f^{*}(\boldsymbol{x})$ represents the current estimate of the optimal true objective. For SBBO algorithms, we employ a Metropolis-Hastings sampling scheme, as described in Algorithm 1. The proposal distribution randomly selects one of the covariates and proposes another value, chosen uniformly at random, which is accepted according to the acceptance probability in step 4 of Algorithm 1. The chosen cooling schedule incrementally increases $H$ from 1 to 10000 in steps of 250.
  2. 2.
    Once the next evaluation point is determined, query the black box function to obtain a noisy observation of the true objective at that location.
  3. 3.
    Append the new pair of data to the dataset and repeat the process.
For each algorithm, we report the best function value found after $t$ evaluations of the true objective, averaged over 10 runs, with one standard deviation divided by the square root of the number of runs as the margin of error. Note that $t$ varies depending on the problem setting.
4.1 Binary Quadratic Problem
Table 1: Average objective function values and margin of error for each algorithm, obtained after 120 iterations for the binary quadratic problem, 500 iterations for the contamination control problem and 300 iterations for the RNA problem. Bold values denote the best value for each problem.
Algorithm Binary Quadratic Problem Contamination Problem RNA Problem
RS $9.86\pm 0.27$ $21.74\pm 0.04$ $-13.74\pm 0.63$
SA $10.75\pm 0.33$ $21.44\pm 0.04$ $-9.21\pm 0.61$
sbbo-BLr $\boldsymbol{11.24}\pm 0.00$ $21.26\pm 0.01$ $\boldsymbol{-22.65}\pm 0.59$
sbbo-BNN $10.90\pm 0.22$ $21.29\pm 0.03$ $-16.08\pm 0.56$
sbbo-GPr $10.76\pm 0.34$ $\boldsymbol{21.20}\pm 0.00$ $-15.05\pm 0.92$
sbbo-NGBdec $10.65\pm 0.28$ $21.38\pm 0.02$ $-13.82\pm 0.59$
sbbo-NGBlinCV $11.10\pm 0.09$ $21.37\pm 0.03$ $-14.04\pm 0.62$
The aim here is to maximize a quadratic function while incorporating $l_{1}$ regularization. This objective function can be expressed as $\boldsymbol{x}^{\top}Q\boldsymbol{x}-\lambda\|\boldsymbol{x}\|_{1}$, where $\boldsymbol{x}$ is a binary vector of dimension $d=10$. The matrix $Q$ is a random $d\times d$ matrix with entries generated as follows: each element is simulated as independent standard Gaussian, and the resulting matrix is then element-wise multiplied by another $d\times d$ matrix $K$. The elements of matrix $K$ are defined as $K_{ij}=\exp\left[-(i-j)^{2}/L_{c}^{2}\right]$, where the correlation length $L_{c}^{2}$ governs how rapidly the values of $K$ decay away from the diagonal. A larger value of $L_{c}^{2}$ results in a denser matrix $Q$, which makes the optimization task more challenging. In our experiments, we set $\lambda$ to 0 and fixed $L_{c}^{2}$ at 10.
Figure 1: BQP problem: average objective function values by number of evaluations, colored by algorithm. Large objective values indicate better performance.
The performance comparison of the different algorithms over 120 iterations is illustrated in Figure 1 and detailed in the second column of Table 1. As can be seen, SBBO outperforms RS and SA for most surrogate models. Notably, SBBO with sparse Bayesian linear regression achieves the highest objective value, closely followed by SBBO with NGB employing a sparse linear model as the base learner. An essential aspect is the assessment of performance with a limited number of objective function evaluations, as evaluating the true objective is the common bottleneck in most BO applications. Remarkably, SBBO-BLr, SBBO-BNN, and SA demonstrate good performance for low number of evaluations ($<$25) (Fig. 1). However, it is worth noting that SA appears to get stuck in local optima around the 65th evaluation and SBBO-BNN performance does not grow as much as that of SBBO-BLr.
In this experiment, with a search space cardinality of 1024, it is feasible to compute the global maximizer of the objective function via complete ennumeration (11.24). Table 2 displays the deviation between the optimal solution discovered within 120 function evaluations for each algorithm and the global optimum. Notably, sbbo-BLr consistently identifies the optimal solution within 120 $\ll$ 1024 iterations across all 10 repetitions of the experiment.
Table 2: Binary quadratic problem: average distance to global optimum achieved after 120 iterations with margin of error. Bold indicates best average performance.
Algorithm Distance to global optimum
RS $1.38\pm 0.23$
SA $0.49\pm 0.33$
sbbo-BLr $\boldsymbol{0.00}\pm 0.00$
sbbo-BNN $0.34\pm 0.22$
sbbo-GPr $0.48\pm 0.34$
sbbo-NGBdec $0.59\pm 0.28$
sbbo-NGBlinCV $0.14\pm 0.09$
4.2 Contamination Control Problem
The contamination control problem was proposed by Hu et al. (2010). Consider a food supply with $d=25$ stages that may be contaminated with pathogenic microorganisms. $Z_{i}$ denotes the fraction of food contaminated at the $i$-th stage of the supply, for $1\leq i\leq d$. $Z$ evolves according to the random process described as follows. At stage $i$, a prevention effort (with fixed cost $c_{i}$) can be made that decreases contamination a random rate $\Gamma_{i}\in(0,1)$. If no prevention is taken, contamination spreads at a random rate $\Lambda_{i}\in(0,1)$. This results in the following recursive equation
$Z_{i}=\Lambda_{i}(1-x_{i})(1-Z_{i-1})+(1-\Gamma_{i}x_{i})Z_{i-1}$
where $x_{i}\in\{0,1\}$ is the decision variable associated with the prevention effort at stage $i$, and $x_{i}=1$ denotes the the prevention effort was taken. The goal is to decide, at each stage $i$, whether to implement a prevention effort. The problem is constrained in that the cost should be minimized while ensuring that at each stage, the fraction of contaminated food does not exceed an upper limit $U_{i}$ with probability at least $1-\epsilon$. The random variables $\Gamma_{i}$ and $\Lambda_{i}$ as well as the initial fraction of contaminated food $Z_{1}$ follow beta distributions. We set $U_{i}=0.1$ for all $i$ and $\epsilon=0.05$, thus adopting the same framework as in Baptista and Poloczek (2018).
We consider the Lagrangian relaxation of the problem that is given by
$\arg\min_{x}\sum_{i=1}^{d}\left[c_{i}x_{i}+\frac{\rho}{T}\sum_{k=1}^{T}\left(1_{\{Z_{ik}>U_{i}\}}-(1-\epsilon)\right)\right]+\lambda\|x\|_{1}$
where each violation is penalized by $\rho=1$. We set $T=100$. An $l_{1}$ regularization term with parameter $\lambda=0.0001$ was added to encourage the prevention efforts to occur at a small number of stages.
Figure 2: Contamination control problem: average objective function values by number of evaluations, colored by algorithm. Small objective values indicate better performance.
The performance comparison of the different algorithms over 500 iterations is illustrated in Figure 2 and detailed in the third column of Table 1. For a small number of function evaluations, simulated annealing performs very well, but quickly gets stuck in a local optimum. SBBO-BLr, on the other hand, showcases consistent performance across both low and high numbers of function evaluations. Notably, SBBO with a Tanimoto Gaussian Process as the surrogate probabilistic model yields the most accurate estimation of the global minimum within 500 function evaluations. However, during initial iterations, this approach falls short of the performance achieved by SBBO-BLr and SA. Importantly, above $\sim 350$ function evaluations, all SBBO-based methods outperform SA and RS.
4.3 RNA design
An RNA sequence can be represented as a string $A=a_{1}\cdots a_{p}$ of length $p$, where each $a_{i}\in\{A,U,G,C\}$ corresponds to a nucleotide letter. The secondary structure of this RNA is defined as the arrangement of base pairs that minimizes its free energy. Several RNA folding algorithms, such as the one proposed by Zuker and Stiegler (1981), employ dynamic programming and a thermodynamic model to estimate the Minimum Free Energy (MFE) of a given sequence. However, the computational complexity of these algorithms is $\mathcal{O}(p^{3})$, which becomes a bottleneck when dealing with a large number of RNA sequences. To address this challenge, we frame the RNA sequence optimization problem as in Dadkhahi et al. (2022): given a sequence of length $p$, the objective is to identify the RNA sequence that forms a secondary structure with the lowest MFE. In our optimization process, we utilize the well-known RNAfold package developed by Lorenz et al. (2011) to assess the MFE for a given RNA sequence. Our ultimate goal is to discover the RNA sequence with the lowest MFE while minimizing the number of calls made to the MFE evaluator. The results that follow correspond to sequences of length $p=30$.
Figure 3: RNA problem: average objective function values by number of evaluations, colored by algorithm. Small objective values of minimum free energy (MFE) indicate better performance.
Figure 3 and the last column of Table 1 showcase the performance comparison of various algorithms over 300 iterations. Notably, Simulated Annealing significantly under-performs, even when compared to Random Local Search, possibly due to the extensive search space (cardinality of $2^{60}$). Remarkably, SBBO-BLr outperforms its counterparts by a significant margin. Although SBBO with a Bayesian Neural Network as a surrogate model excels in scenarios with a low number of function evaluations and finds the second-best maximizer after 300 objective function evaluations, its performance remains notably distant from that of SBBO-BLr.
4.4 General Comments
While most Bayesian Optimization literature have utilized Gaussian Processes as surrogate probabilistic models, our empirical study suggests that, for combinatorial optimization problems, alternative Bayesian models may lead to superior performance. Notably, in two out of three experiments, Sparse Bayesian linear regression with pairwise interactions demonstrated clear superiority over GPs and maintained competitiveness in the remaining experiment. This underscores the significance of exploring a broader range of Bayesian models within the realm of Bayesian Optimization.
Given that many Bayesian models lack closed-form expressions for posterior predictive distributions, our study emphasizes the relevance of developing algorithms to optimize acquisition functions in Bayesian Optimization when only sampling access to posteriors is available. This is particularly important as inference in most Bayesian models relies on simulation. SBBO aligns precisely with this objective.
5 Conclusion and Future Work
We have proposed SBBO, a simulation-based framework for optimizing acquisition functions in Bayesian Optimization that just requires sampling access to posterior predictive distributions of surrogate models. This flexibility enables the utilization of any Bayesian model whose posterior inference relies on MCMC as surrogate probabilistic model in BO. This can have advantages in non-standard black-box function optimization settings, as in the case of combinatorial search spaces. We have demonstrated empirically the effectiveness of SBBO under several surrogate probabilistic models in three combinatorial optimization problems.
Several future lines of research are possible. From a methodological perspective, we have only considered the greedy one-step look-ahead BO. SBBO could be extended to multi-step look-ahead BO. However, this extension is not direct. From a computational perspective we would suffer from the standard inefficiencies of dynamic programming. Some ideas in Bayesian sequential design might be relevant Wathen and Christen (2006); Brockwell and Kadane (2003); Drovandi et al. (2014). In terms of sampling strategies, we have proposed Metropolis-Hastings and Gibbs samplers to sample from the augmented distribution. Other options could be adopted to better explore high-dimensional multimodal surfaces, such as particle methods Amzal et al. (2006). In addition, when dealing with continuous search spaces, MCMC samplers that leverage gradient information, such as Hamiltonian Monte Carlo or the Metropolis-adjusted Langevin algorithm could be used.
On the applications side, given that Bayesian linear regression with pairwise interactions seems to work well, a model that captures interactions between three or more covariates is of interest. However, explicitly accounting for beyond-pairwise interactions becomes unfeasible when the dimensionality of the search space increases. As an alternative, some non-parametric approaches such as Bayesian additive regression tress Chipman et al. (2010) that capture relevant interactions between more than two covariates could be utilized. Finally, SBBO could be used with likelihood-free surrogate models, from which posterior predictive samples could be obtained via Approximate Bayesian Computation.
Acknowledgements: The authors gratefully acknowledge Spain’s consul in Chicago for his valuable help facilitating this collaboration during the pandemic. This project is based upon work supported by the Spanish state investigation agency under the Proyectos de Generación de Conocimiento 2022 grant No. PID2022-137331OB-C33.
Disclosure statement: The authors report there are no competing interests to declare.
SUPPLEMENTARY MATERIAL
Appendix A Empirical assessment of convergence
We include Figure 4 to provide empirical guarantees of SBBO convergence. To make this plot, we initialized the contamination problem described in Section 4.2 with a dataset $\mathcal{D}_{1:100}$ of 100 of pairs of covariates and outcomes. Metropolis Hastings SBBO (Algorithm 1) with Tanimoto Gaussian Process regression as surrogate probabilistic model was used to find the subsequent evaluation location $x_{101}$. This process was repeated 40 times to account for the stochasticity of the SBBO algorithm. In each repetition, the cooling schedule incrementally increases $H$ from 50 to 2500 in steps of 10. For each $H$, we approximate the posterior predictive utility of the solution found by SBBO for that value of $H$ via Monte Carlo. Expected improvement was used as utility function. $H$ is represented in the x axis and the corresponding posterior predictive utility in the y axis. Each gray line corresponds to a repetition and the red line indicates the average.
Figure 4: SBBO with Metropolis-Hastings convergence